OpenStudy (anonymous):

Write 6+i / 6-i in a + bi form

OpenStudy (anonymous):

what would happen if you multiplied the entire thing by \[\frac{6+i}{6+i}\]

OpenStudy (anonymous):

I would get 36-1/ 36-1? I'm not sure for the bottom half

OpenStudy (anonymous):

(6+i)^2 is not equal to 36-1 try that again

OpenStudy (anonymous):

anyways the bottom will become a real number we use difference of squares to get rid of the complex number on the bottom once you do this, you will have a complex number over a real number so distribute the real number and you will get your answer

OpenStudy (anonymous):

1 + 12i?

OpenStudy (anonymous):

is that your answer? i dont actually do these problems, i usually just give steps

OpenStudy (anonymous):

I guess it is. I'm really lost though..

OpenStudy (anonymous):

ok just looking at the numerator (6+i)(6+i) this is the same as 6(6+i)+i(6+i)= what is the answer to that? please write down your steps

OpenStudy (anonymous):

6(6+i)+i(6+i)= 36+6i + 6i+i^2 = 36-1+12i you go the 12i correct just not other part so the numerator then becomes 35+12i

OpenStudy (anonymous):

now you already did the denominator (6+i)(6-i)=36-1=35

OpenStudy (anonymous):

so what you are left with is \[\frac{35+12i}{35}\] can you finish this now?

OpenStudy (anonymous):

Oh that is what I got. I assumed that 35/35 was 1 but Im pretty sure I got it know thankyou so much.

OpenStudy (anonymous):

well you need to change it to a+bi form so \[\frac{35+12i}{35}=\frac{35}{35}+\frac{12i}{35}=1+\frac{12}{35}i\]

OpenStudy (anonymous):

Oh I see my mistake, I am supposed to add the denomenator under 12i as well. Thankyou.

OpenStudy (phi):

The main idea is multiply top and bottom by the "complex conjugate" of the denominator if you have a+bi, the complex conjugate is a-bi (a+bi)(a-bi)= a^2 - (bi)^2 = a^2 -i^2 b^2 i^2= -1 , so you get a^2+b^2 which is a real number for the denominator